# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ 4.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures. Below, we discuss some of these major threats.
  • Autonomy abuse. The foremost threat is the autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs . Unlike deterministic models, these Agentic AI systems dynamically generate actions, complicating efforts to define and enforce safe operational states .
  • Persistent memory. Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential prompt injections. Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms .
  • Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [53] and AutoGen [98] . These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
  • Goal Misalignment and Exposure. Agentic AI systems often operate with internalized or user-specified goals. Improper scoping or exposure of goal representations can lead to prompt extraction, misalignment, or competitive agent behaviors, particularly when third-party tools or APIs are invoked dynamically during task completion.
  • Tool Misuse and External API Exploits. Agents capable of invoking external tools (e.g., via APIs, code execution, or web access) present novel attack surfaces. Misuse of these tools, intentional or accidental, can trigger financial costs, Distributed Denial of Service (DDoS) behavior, or legal violations. Sandboxing and invocation tracing remain limited in most open-source frameworks.
  • Multi-Agent Collusion or Drift. Emergent coordination among agents, while useful for collaboration, can also lead to collusion, competitive drift, or polarization of outputs, especially in self-reinforcing memory loops or agent swarms. Such behavior complicates interpretability and may undermine safety alignment goals.
4.2 Taxonomy of Risks
We propose a taxonomy of risks associated with AMAS, categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 4. These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS. Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
  • Adversarial Attacks. Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps. These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [99]. For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [30]. Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
  • Data Leakage. The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information. In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [100]. For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls. Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
  • Agent Collusion and Mode Collapse. Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [101]. This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies. For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance. Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
  • Emergent Behaviors. The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [102]. Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability. Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [103], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals . These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
@@ 4.2 Taxonomy of Risks
Several real-world examples shows the consequences risks in deployed or experimental Agentic AI systems.
  • Prompt Leakage in Agentic AI Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [104], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [105]. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks.
  • Collusive Failure in ChatDev. In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of a faulty design [86]. Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations. This scenario highlights the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures [37].
  • Simulation Attack in Swarm Robotics. In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion [106]. This incident demonstrates the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences. The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior [107].
  • Memory Poisoning in Multi-Agent Chatbots. In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer. This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones. This incident highlights the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities [108].
  • System Prompt Drift in Autonomous Memory Agents. In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated. This led to hallucinated goals and emergent behaviors misaligned with initial intentions [109].
Next, we discuss the AI TRiSM framework and map its core principles to a AMAS and how it can mitigate these aforementioned risks.
5 The TRiSM Framework for Agentic Multi-Agent Systems
5.1 Overview of AI TRiSM and Its Core Components
The TRiSM Framework, short for Trust, Risk, and Security Management, offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors. In the context of Agentic AI, where LLMs are deployed as self-directed agents capable of perception, reasoning, and interaction, the TRiSM framework provides a critical foundation for ensuring robustness, transparency, and responsible deployment.
@@ 5.1 Overview of AI TRiSM and Its Core Components
  • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness. Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [54]. This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
@@
  • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies. This includes technical risks such as algorithmic bias [41], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [11]. Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
@@
  • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access. This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [38]. Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
Originally highlighted in industry guidelines for AI governance [110], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems. The TRiSM framework addresses trust concerns in Agentic AI by focusing on key pillars, which are: Explainability, ModelOps, Application Security, Model Privacy and Goverance [64, 13]. Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
Figure 5: TRiSM-aligned architecture for Agentic AI systems. The system is organized into four layers: Trust and Explainability (e.g., reasoning traces, user feedback), ModelOps (prompt versioning, CI/CD, orchestration), Planning and Reasoning (multi-agent CoT), and Model Privacy (encrypted memory, differential privacy, audit logs). A cross-cutting Governance Plane enforces oversight via policies and logging. External tools are accessed through a secure API gateway. The user interface connects back to the trust and governance layers for human-in-the-loop (HITL) oversight. Abbreviations used: TRiSM = Trust, Risk, and Security Management; CoT = Chain-of-Thought; CI/CD = Continuous Integration and Deployment; DP = Differential Privacy.
Table 4: TRiSM pillars for Agentic (LLM-based) multi-agent systems: core controls, techniques, risks, evaluation facets, example systems, and governing frameworks.
Pillar Core controls (keywords) Techniques / patterns / artifacts Key risks mitigated Evaluation facets & example metrics Example systems / patterns Primary standards / frameworks
@@
Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [111]; LIME/SHAP [112, 113]; decision-provenance graphs [114, 115]; multi-agent SHAP [116]; attention maps [117, 118]; role-based interpretability [119]; RAG-linked justifications [120]; prompt/action logs [121, 122] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [123] SciAgent [120]; MetaGPT [124]; ChatDev [30]; AutoGen [57] EU AI Act Arts 13–14 [125]; NIST AI RMF / GAI profile [3, 126]; OECD AI [127]; ISO/IEC 24029-1 [32]
@@
ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [128] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent (hierarchical monitoring) [128] ISO/IEC 42001 (AIMS) [ISO_IEC_42001_2023]; ISO/IEC 42005 [129]; ISO/IEC 23894 [iso23894 or [130]]; NIST AI RMF / GAI profile [3, 126]; ModelOps [131, 132]
@@
Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [133, 58]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [134]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [135]; enterprise dashboards (override/stop) [135]; ChemCrow safety pauses [136] OWASP LLM Top-10 [137, 138]; OECD Robustness [127]
@@
Model Privacy & Data Protection DP; anonymization/pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [139]; k-anonymity [140]; MPC/PUMA [puma2024 or [141]]; HE/FHE pipelines [142]; TEEs [143]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [34]; CCPA/CPRA [144]; HIPAA §164 [35]; ISO/IEC 27001/27701 [iso27001, iso27701]
@@
Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [125]; NIST AI RMF [3]; OECD AI [127]; ISO/IEC 42001/42005 [align keys]; ISO/IEC 23894 [130]
@@ 5.2 Mapping Pillars of TRiSM with Agentic AI
In this section, we examine how the pillars of TRiSM align with Agentic AI (or AMAS) and summarize our syntheis in Table LABEL:tab:trism-pillars. For each pillar, we outline its relevance and implications within the context of Agentic AI.

### v5

@@ 3.1 Threats in Agentic AI
Agentic AI systems introduce security and reliability concerns beyond single-agent LLM pipelines:
  • Autonomy abuse. Agents with delegated authority may misinterpret objectives or execute harmful plans due to reasoning errors or manipulated inputs.
  • Persistent memory. Long-term memory improves continuity but creates a contamination channel: injected or low-quality content can persist and propagate through shared stores without strong versioning and audit controls.
  • Agent orchestration. Centralized or distributed orchestrators mediate roles and workflows; compromise or misconfiguration can misroute information and trigger cascading failures.
  • Goal misalignment and exposure. Poorly scoped or exposed goal representations can enable prompt extraction, misalignment, or adversarial steering, especially when third-party tools are invoked.
  • Tool misuse and external API exploits. Tool-enabled agents introduce new attack surfaces (e.g., unsafe code execution, costly actions, policy violations); sandboxing and invocation tracing are often inconsistent across open-source frameworks.
  • Multi-agent collusion or drift. Agents can converge on mutually reinforcing errors (groupthink/mode collapse) or drift in self-reinforcing loops, complicating interpretability and safety alignment.
3.2 Risks Taxonomy in AMAS
We group AMAS risks into four classes—adversarial attacks, data leakage, agent collusion, and emergent behaviors—as shown in Fig. 10.
Figure 10: Overview of key risks and threats in agentic AI systems. The left panel outlines system-level threats (e.g., autonomy abuse, persistent memory, orchestration). The right panel summarizes emergent AMAS risk classes arising from multi-agent interaction, memory persistence, and decentralized decision-making.
  • Adversarial attacks. AMAS are vulnerable to prompt injection, engineered reasoning traps, and other manipulations that can cascade across agents; e.g., role-swapping attacks in ChatDev exploit inter-agent dependencies [miao2025autonomous, qian2023chatdev].
  • Data leakage. Shared memory and inter-agent communication increase the chance of unintended disclosure of sensitive information, particularly when boundaries, sanitization, or access controls are weak [tran2025multi].
  • Agent collusion and mode collapse. Collaboration mechanisms can produce groupthink, where agents reinforce the same flawed assumptions and converge on suboptimal outputs [feng2025multi].
  • Emergent behaviors. Complex interactions among agents, tools, and memory can yield unpredictable strategies that evade conventional testing (e.g., shortcutting verification), motivating continuous monitoring and adaptive controls [dawid2025agentic].
@@ 3.2 Risks Taxonomy in AMAS
The following examples illustrate how these risks can appear in practice:
  • Prompt leakage. In AutoGPT-style setups, recursive prompt augmentation and weak memory controls can expose sensitive content when stored tokens later surface in summaries or logs [Euler2023AutoGPTRCE, invicti_prompt_injection_ebook].
  • Collusive failure. In ChatDev-like workflows, shared planning errors can propagate when agents validate each other without external checks or objective feedback [qian2024chatdevcommunicativeagentssoftware].
  • Simulation attack in swarms. Misleading environment assumptions can induce coordination failures (e.g., congestion and task incompletion) in multi-agent planning settings [StrDorFri2024, ligot2022using].
  • Memory poisoning. Feedback buffers can be polluted (e.g., sarcastic or adversarial signals) and later used for policy updates, degrading behavior without validation gates [atkins2023memorypoisoning].
  • System prompt drift. Self-appended context without versioning can shift prompts over time and induce hallucinated goals or misaligned behaviors [liu2023lostmiddlelanguagemodels].
Key Insights
We identified AMAS-specific risks that arise from autonomy, shared memory, orchestration, and tool use, and organized them into four emergent classes (Fig. 10). Across the taxonomy, failures can cascade across agents and persist through memory substrates, while emergent behaviors may evade conventional test suites. These observations motivate defense-in-depth controls spanning prompt hygiene, memory sanitization, orchestrator authentication, tool sandboxing, and continuous monitoring.
4 TRiSM Framework for Agentic Multi-Agent Systems
4.1 Overview of AI TRiSM
The TRiSM Framework offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors. Below, we discuss core components of TRiSM framework.
@@ 4.1 Overview of AI TRiSM
  • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness. Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [borghoff2025human]. This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
@@
  • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies. This includes technical risks such as algorithmic bias [raza2024exploring], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [fang2025trustworthyaisafetybias]. Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
@@
  • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access. This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [raza2025justhumansneedvaccines]. Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
Table 4: TRiSM pillars (Part 1): Core controls, techniques, and key risks mitigated.
Pillar Core Controls Techniques / Patterns Key Risks Mitigated
@@
Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [sanwal2025layered]; LIME/SHAP [ribeiro2016should, lundberg2017unified]; decision-provenance graphs [tupayachi2024towards, bai2024derived]; multi-agent SHAP [tian2025outlook]; attention maps [liu2024ava, hu2024overview]; role-based interpretability [tomsett2018interpretable]; RAG-linked justifications [ghafarollahi2024sciagents]; prompt/action logs [gu2024artificial, chin2025automating] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration
@@
ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [wang2025megaagent] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps
@@
Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [beurer2025design, qiao2024agent]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [standen2025adversarial]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks
@@
Privacy & Data Protection DP; anonymization; pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [dwork2006calibrating]; k-anonymity [sweeney2002k]; MPC/PUMA [lindell2009secure]; HE/FHE pipelines [gentry2009fully]; TEEs [costan2016intel]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention
@@
Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions
Table 5: TRiSM pillars (Part 2): Evaluation facets, example systems, and governing standards.
Pillar Evaluation Facets & Metrics Example Systems Standards / Frameworks
Explainability / Trust Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [raza_fair_2024] SciAgent [ghafarollahi2024sciagents]; MetaGPT [ke2025meta]; ChatDev [qian2023chatdev]; AutoGen [wu2023autogen] EU AI Act Arts 13–14 [eu_ai_act]; NIST AI RMF / GAI profile [ai2023artificial, ai2024artificial]; OECD AI [oecd_ai_catalogue]; ISO/IEC 24029-1 [iso24029-1]
ModelOps (Lifecycle) Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent [wang2025megaagent] ISO/IEC 42001 (AIMS); ISO/IEC 42005 [ISO_IEC_42005_2025]; ISO/IEC 23894 [ISOIEC23894_2023]; NIST AI RMF [ai2023artificial, ai2024artificial]; ModelOps [lefevre2022modelops, Sinha2025ModelOps]
Application Security Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [chen2025human]; enterprise dashboards [chen2025human]; ChemCrow safety pauses [bran2023chemcrow] OWASP LLM Top-10 [owasp2024llmtop10, owasp2025agentic]; OECD Robustness [oecd_ai_catalogue]
Privacy & Data Protection Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [gdpr25]; CCPA/CPRA [oag_ccpa_2024]; HIPAA §164 [hipaa164]; ISO/IEC 27001/27701
Governance User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [eu_ai_act]; NIST AI RMF [ai2023artificial]; OECD AI [oecd_ai_catalogue]; ISO/IEC 42001/42005; ISO/IEC 23894 [ISOIEC23894_2023]
@@ 4.2 Mapping Pillars of TRiSM with Agentic AI
Lifecycle Governance Plane (TRiSM Pillar) policies $\cdot$ accountability $\cdot$ compliance $\cdot$ oversight $\cdot$ traceability
Trust & Explainability (TRiSM Pillar) reasoning/provenance traces $\cdot$ explainer agent $\cdot$ user feedback $\cdot$ audit logs ModelOps / Lifecycle Management (TRiSM Pillar) prompt+agent config versioning $\cdot$ CI/CD safety gates $\cdot$ monitoring/rollback $\cdot$ incident workflow Application Security (TRiSM Pillar) prompt hygiene $\cdot$ allowlists $\cdot$ tool sandboxing $\cdot$ least privilege (RBAC) $\cdot$ anomaly detection Model Privacy & Data Protection (TRiSM Pillar) encrypted/scoped memory $\cdot$ data minimization $\cdot$ DP where applicable $\cdot$ access logs $\cdot$ retention/consent Planning & Reasoning (Agent Core Layer; non-TRiSM) multi-agent planning $\cdot$ task decomposition $\cdot$ CoT/ToT/ReAct $\cdot$ inter-agent coordination
User / HITL UI
confirm $\cdot$ override $\cdot$ stop
feedback loop
Secure API Gateway
authN/authZ $\cdot$ egress control
tool isolation $\cdot$ policy checks
External Tools
web $\cdot$ DB $\cdot$ code
enterprise APIs
Figure 11: TRiSM-aligned architecture for Agentic AI systems. The five TRiSM pillars are shown explicitly and in the same order as defined in the framework: Trust & Explainability, ModelOps, Application Security, Model Privacy & Data Protection, with a cross-cutting Lifecycle Governance plane. Planning & Reasoning is included as the agent-core execution layer (non-TRiSM).
Originally highlighted in industry guidelines for AI governance [litan2024ai], TRiSM also provides a structured approach to managing the unique challenges of AMAS. In this work, we focus on five key TRiSM pillars: Explainability, ModelOps, Application Security, Model Privacy, and Governance [kaur2024building, habbal2024artificial]. These five pillars also reflect established risk-surface decompositions in the NIST AI RMF [ai2023artificial], ISO/IEC 42001 [iso42001], and the EU AI Act [eu_ai_act], where explainability, lifecycle management, security, privacy, and governance are treated as distinct and non-overlapping risk domains. We summarize this mapping in Tables LABEL:tab:trism-pillars-1 and LABEL:tab:trism-pillars-2. Figure 11 illustrates how these five TRiSM pillars map onto the core architectural layers of an LLM-based agentic AI system.
Next, we discuss these pillars and how they can help in AMAS.